A life insurance salesman sells on the average 3 life insurance policies per week, assuming Poisson distribution is followed. Compute the probability he sells more than 2 or more policies for duration of 2 weeks.



1
Expert's answer
2021-05-07T14:23:13-0400

We have given that,

"\mu = 3"

We have to find "P(X \ge 2)"

Using the Poisson distribution,

"P(X) = \dfrac{e^{-\mu}\mu^x}{ x!}"

then,

"P(X \ge 2) = 1-(P(0)+P(1))"


"= 1- \dfrac{e^{-3}}{1} - \dfrac{e^{-3}(3)}{1}"


"= 0.811"